# GPU-Accelerated Parallel Selected Inversion for Structured Matrices Using sTiles

### v1

@@ III-C Two-Phase Algorithm for Selected Inversion
The Directed Acyclic Graphs (DAGs) in Figure 4 directly represent the task distribution and parallelism strategies implemented in the two-phase algorithm detailed above. The graphs demonstrate how tasks are assigned to different cores, with each color representing the tasks handled by a specific core. While this visualization uses matrices of size $6\times 6$ tiles, the distribution and parallelism would become even more apparent for larger matrices, where the increased number of tasks leads to a more complex workload distribution.
III-D GPU Acceleration for Parallel Selected Inversion
@@
While our current implementation assumes that the selected tiles fit in GPU memory, the algorithm could be extended to support out-of-core execution for larger matrices that exceed GPU capacity. This would involve overlapping data movement and computation while maintaining efficient static scheduling. Such strategies have been successfully applied in the context of out-of-core Cholesky factorization on modern GPU architectures, as demonstrated in [17]. Incorporating similar techniques into sTiles would enable scalable selected inversion on next-generation heterogeneous systems such as the NVIDIA Grace Hopper Superchip with a unified memory subsystem on the CPU host.
@@ IV-A Experimental Setup and Software
Our computational experiments were conducted on two high-performance computing (HPC) systems, each designed to assess the performance of sTiles under different computational paradigms:
  • CPU Server: A dual-socket 26-core system featuring Intel® Xeon® Gold 6230R processors, each equipped with 52 cores total operating at 2.10 GHz, with a total L3 cache of 71.5 MB.
  • GPU Node: A dedicated compute node incorporating 64 AMD EPYC 7713 CPU cores running at 1.99 GHz, complemented by a single NVIDIA A100-SXM4 GPU, which operates at 1.16 GHz and includes 80 GB of high-bandwidth memory (HBM2).
This setup facilitates a comprehensive evaluation of sTiles across CPU-only and GPU-accelerated configurations, providing insights into its computational efficiency and scalability.
To benchmark our parallel selected inversion algorithm, we compare it against Panua-PARDISO 8.2, a state-of-the-art direct solver for sparse linear systems utilizing $LL^{T}$ factorization, optimized for shared-memory parallelism. This solver employs advanced numerical techniques to ensure efficient factorization and inversion, making it a robust reference for our comparative analysis.
Additionally, SelInv 2.0.0 was initially considered for inclusion in our performance assessment. However, several issues precluded a fair and meaningful comparison: (1) the software is no longer actively maintained, (2) the only functional configuration we identified runs with a single MPI process and mandates a specific ordering option, and (3) certain ordering configurations fail to work. Moreover, SelInv 2.0.0 relies on SuperLU for its inverse computation backend, whereas our approach assumes $LL^{T}$ factorization. Due to these constraints, we excluded SelInv 2.0.0 from our evaluation.
@@ IV-B Performance Evaluation
TABLE I: Matrix properties used in Cholesky factorization and selected inversion experiments for sTiles – Set 1. These matrices reflect the arrowhead structures that commonly arise in INLA-based models [4].
ID Size Bandwidth Arrowhead Thickness Density (%)
@@
Table I summarizes the structured matrices utilized in our evaluation, encompassing variations in size, bandwidth, and sparsity levels. These matrices are carefully chosen to represent practical applications, such as statistical models with structured sparsity patterns. Specifically, the arrowhead thickness corresponds to the number of fixed effects in statistical models, with values ranging from 10 (moderate case) to 200 (extreme case). The benchmarking aims to assess the computational efficiency of sTiles in handling these structures and to compare its performance against Panua-PARDISO 8.2, a state-of-the-art solver.
@@
    Factorization and Selected Inversion: Cholesky factorization is performed, followed by selected inversion using sTiles. The execution time is compared against Panua-PARDISO 8.2.
@@
    Performance Evaluation: Execution times are recorded across increasing core counts (1, 2, 4, 8, 16, 32, and 52 cores, the maximum available on our test system). The best execution time for each configuration is selected.
@@
    Performance Visualization: Results are analyzed in terms of absolute runtime and relative speedup compared to Panua-PARDISO.
@@
Figure 5 demonstrates the performance advantage of sTiles over Panua-PARDISO across different matrix sizes. The performance gap becomes more pronounced as matrix size increases (IDs 7-18), with sTiles leveraging dense tiles to mitigate computational overhead and memory bandwidth limitations. In some cases, sTiles achieves up to 13.49× speedup compared to Panua-PARDISO.
IV-C Scalability Evaluation
Figure 6: Execution times of sTiles and PARDISO for matrices of size $10K$ (left) and $500K$ (right) across different core counts. The y-axis is in logarithmic scale.
Figure 6 presents the execution times of sTiles and Panua-PARDISO across different core counts for two representative matrix sizes: $10K$ and $500K$. For the smaller matrix, both solvers exhibit strong scalability, maintaining comparable performance across all configurations. However, for the larger matrix, sTiles demonstrates superior scalability, achieving consistently lower execution times as the core count increases. This performance advantage is attributed to sTiles’ efficient exploitation of parallelism.
@@ IV-D Impact of Sparsity on Performance
The impact of sparsity on the computational performance of selected inversion is analyzed using a set of structured matrices with varying densities. Table II summarizes the properties of these matrices, including size, bandwidth, arrowhead thickness, and density. The density values, which exclude the arrowhead portion, range from 0.010% to 4.101%, providing a broad spectrum of sparsity levels for evaluation.
Figure 7 presents the inverse computation times for both PARDISO and sTiles solvers across matrices with increasing density. The results illustrate a clear distinction in solver behavior depending on the sparsity characteristics. For matrices with very low density (below 0.1%), PARDISO exhibits a faster performance than sTiles, primarily due to its optimized multifrontal structure for handling highly sparse matrices. However, as the density increases beyond this threshold, the performance of sTiles stabilizes, while PARDISO experiences significant computational overhead.
The results indicate that sTiles maintains consistent computational times across a wide range of density values, demonstrating its robustness for handling moderately sparse to dense matrices. In contrast, PARDISO’s runtime increases substantially with density, reflecting the growing complexity of fill-in and symbolic factorization in direct solvers. For high-density matrices (greater than 1%), sTiles outperforms PARDISO, making it a more scalable approach for dense and structured sparse matrices.
@@
TABLE II: Matrix properties used in Cholesky factorization and selected inversion experiments for sTiles – Set 2. Density values exclude the arrowhead part.
@@
Figure 7: Selected inverse computation times of sTiles and PARDISO across varying matrix densities. The top plot illustrates results for matrices with a bandwidth of 1500, while the bottom plot corresponds to matrices with a bandwidth of 3000. The x-axis represents matrix IDs, with density values displayed beneath each ID, while the y-axis uses a logarithmic scale to capture the variations in computation time.
IV-E Full Matrix Inversion Performance
This section examines the performance of sTiles in full matrix inversion and compares it against PLASMA [16, 15], highlighting the impact of scheduling strategies. PLASMA employs dynamic scheduling, while sTiles uses static scheduling to increase data locality and minimize scheduling overhead.
Figures 8 and 9 present the execution times of both libraries across different tile sizes and core counts for matrix ID 5 from Table I. The main observations are:
  • sTiles consistently outperforms PLASMA, achieving lower execution times across all configurations.
  • PLASMA suffers from performance degradation beyond 16 cores, particularly for small tile sizes (40 and 80). This is attributed to dynamic scheduling overhead and thread contention.
  • sTiles scales efficiently, maintaining a steady decrease in execution time up to 52 cores, demonstrating better parallel efficiency.
  • Tile size sensitivity is lower in sTiles, meaning it achieves strong performance across different tile sizes without requiring extensive tuning.
@@ IV-E Full Matrix Inversion Performance
These results indicate that sTiles remains efficient even in full matrix inversion scenarios, maintaining strong scalability and robust performance across all tile sizes, while PLASMA’s dynamic scheduling leads to a large time complexity at high core counts. To further emphasize this, we plot the full inversion comparison using the best tile size for each library (PLASMA: 280, sTiles: 320) in Figure 10.
Figure 10: Full inversion comparison using the best tile size for each library (PLASMA: 280, sTiles: 320). sTiles consistently achieves lower execution times across all core counts.
@@ IV-F Accelerating Selected Inversion on GPU
To evaluate the benefits of GPU acceleration, we performed selected inversion on a large matrix of size 50,010, with a bandwidth of 15,000, an arrowhead thickness of 10, and a density of 0.3123. This configuration represents a computationally intensive problem well-suited for GPU execution, as the large bandwidth leads to dense blocks and a high ratio of floating-point operations to memory accesses.
Following the experimental setup in [14], we used a tile size of 120 for CPU computations and a larger tile size of 600 for GPU computations. While the GPU hardware itself offers inherent performance advantages, additional gains are achieved through preprocessing optimizations during factorization. Notably, differences in tile size influence the structure of the tiled matrix, symbolic factorization complexity, and the distribution of computational tasks across resources.
TABLE III: CPU vs GPU performance comparison for the two-phase selected inversion.
Cores/Streams CPU Time (s) GPU Time (s)
1 533.062 7.044
2 282.714 4.524
4 142.883 4.366
8 75.225 3.954
16 39.575 4.111
32 26.861 3.562
64 20.382 3.949
All GPU results were obtained using the NVIDIA A100 GPU available on the GPU node described in Section IV-A, while CPU results were measured on the dual-socket 26-core Intel Xeon server.
As shown in Table III, the GPU achieved a runtime of 3.949 seconds, compared to 20.382 seconds on 64 CPU cores, yielding a speedup of 5.12$\times$.
These results highlight the efficiency of GPU acceleration for large-scale structured matrix problems, particularly when combined with tile-aware design choices that exploit the high arithmetic intensity and parallelism available on modern GPU architectures.
The observed GPU speedup aligns well with expectations derived from the Roofline Performance Model, which suggests a theoretical peak performance ratio of up to 20$\times$ between CPU and GPU (1 TFLOP/s vs. 20 TFLOP/s). In practice, however, the speedup is moderated by host-to-device data transfer overheads and the inherent limitations of workload concurrency. While the GPU architecture is highly effective for throughput-oriented parallel computations, its ability to reach peak performance depends on the volume and granularity of the workload. When concurrency is limited or data movement becomes a bottleneck, the achievable speedup is reduced, although it may still indicate performance superiority.

### v2

@@ III-D Two-Phase Algorithm for Selected Inversion
The Directed Acyclic Graphs (DAGs) in Figure 5 directly represent the task distribution and parallelism strategies implemented in the two-phase algorithm detailed above. The graphs clearly demonstrate how tasks are assigned to different cores, with each color representing the tasks handled by a specific core. While this visualization uses a matrix of size 6x6 tiles for illustrative clarity, the principles of static task distribution and parallelism it demonstrates are fundamental to the algorithm’s scalability for larger problems. For larger matrices, this structured partitioning leads to a significantly greater number of concurrent tasks, fully exploiting the available computational resources.
(a) Full matrix inversion using 2 cores.
(b) Full matrix inversion using 4 cores.
(c) Arrowhead matrix inversion using 2 cores.
(d) Arrowhead matrix inversion using 4 cores.
Figure 5: Directed acyclic graphs (DAGs) illustrating the task distribution for selected inversion. The top two rows show the DAGs for a full matrix with 2 and 4 cores, respectively; their width illustrates the high degree of potential parallelism. The bottom row shows the significantly more compact DAGs for a structured arrowhead matrix, highlighting the drastic reduction in the number of tasks. Each color represents tasks assigned to a specific core.
III-E Algorithmic Complexity
A matrix of size $n\times n$ is tiled into $N\times N$ blocks of size $b\times b$, so n=Nbn=N\,b. The cost of a single tile operation, such as GEMM or TRSM, is $O(b^{3})$ floating-point operations.
For a full, dense matrix, the complexity of our block inversion algorithm is dominated by GEMM operations, resulting in a total workload of $O(n^{3})$, which is consistent with standard methods.
In contrast, for the selected inversion of an arrowhead matrix with a band of width consisting of $B$ tile blocks, the workload is significantly reduced. By only performing computations that contribute to the selected arrowhead pattern, the complexity becomes:
$W_{\text{selected}}=O(B^{2}nb^{2})$
This complexity is asymptotically lower than the dense $O(n^{3})$ case, making the method highly efficient when the band $B$ is small relative to the number of tiles $N$. Table I illustrates this drastic reduction in the number of required block operations. A detailed derivation of these complexity results is provided in Appendix A.
TABLE I: Comparison of block operation counts for full dense vs. selected inversion.
Inversion GEMM Ops Asymptotic
Type ($N=6$) Trend
Full ($B=6$) 70 $O(N^{3})\to O(n^{3})$
Selected ($B=1$) 10 $O(N)\to O(nb^{2})$
Selected ($B=2$) 26 $O(N)\to O(nb^{2})$
III-F GPU Acceleration for Parallel Selected Inversion
@@
While our current implementation assumes that the selected tiles fit in GPU memory, the algorithm could be extended to support out-of-core execution for larger matrices that exceed GPU capacity. This would involve overlapping data movement and computation while maintaining efficient static scheduling. Such strategies have been successfully applied in the context of out-of-core Cholesky factorization on modern GPU architectures, as demonstrated in [20]. Incorporating similar techniques into sTiles would enable selected inversion to scale to larger problem sizes on next-generation heterogeneous systems such as the NVIDIA Grace Hopper Superchip with a unified memory subsystem on the CPU host.
@@ IV Performance Evaluation and Experimental Results
Figure 6: Visual illustration of matrix structure showing bandwidth and arrowhead thickness. These dimensions influence the density and shape of the matrices used in our performance evaluation. In typical Bayesian inference modeling such as the INLA method [5], the arrowhead thickness corresponds to the number of fixed effects included in the model, while the bandwidth reflects the local structure of random effects.
@@ IV-A Experimental Setup and Software
Our computational experiments were conducted on two high-performance computing (HPC) systems, each used for a specific set of evaluations involving sTiles:
  • CPU Server: A dual-socket 26-core system featuring Intel® Xeon® Gold 6230R processors, with 52 cores total operating at 2.10 GHz and a total L3 cache of 71.5 MB. This system was used for all CPU-only experiments, including scalability, density sensitivity, and full matrix inversion. Additionally, the Panua-PARDISO license is available on this system, enabling direct solver comparisons under a fully licensed environment.
  • GPU Node: A dedicated compute node incorporating 64 AMD EPYC 7713 CPU cores running at 1.99 GHz, complemented by a single NVIDIA A100-SXM4 GPU operating at 1.16 GHz and equipped with 80 GB of high-bandwidth memory (HBM2). This system was used for all GPU-accelerated experiments.
In comparisons between CPU and GPU performance (e.g., Table IV), the CPU baseline was measured using the AMD EPYC 7713 CPU on the GPU node to ensure consistency in hardware and avoid bias introduced by architectural differences.
To benchmark our parallel selected inversion algorithm, we compare it against Panua-PARDISO 8.2, a state-of-the-art direct solver for sparse linear systems utilizing $LL^{T}$ factorization, optimized for shared-memory parallelism. This solver employs advanced numerical techniques to ensure efficient factorization and inversion, making it a robust reference for our comparative analysis. To isolate the impact of the selected inversion algorithm, we compare only the selected inversion time in our performance evaluation. The Cholesky factorization time is excluded from these comparisons. A detailed analysis of the Cholesky phase has already been presented separately in [13].
@@ IV-B Performance Evaluation
TABLE II: Matrix properties used in Cholesky factorization and selected inversion experiments for sTiles – Set 1. These matrices reflect the arrowhead structures that commonly arise in INLA-based models [5].
Matrix ID Size Bandwidth Arrowhead Thickness Density (%)
@@
Table II summarizes the structured matrices utilized in our evaluation, encompassing variations in size, bandwidth, and sparsity levels. These matrices are carefully chosen to represent practical applications, such as statistical models with structured sparsity patterns. Specifically, the arrowhead thickness corresponds to the number of fixed effects in statistical models, with values ranging from 10 (moderate case) to 200 (extreme case), see Figure 6. The benchmarking aims to assess the computational efficiency of sTiles in handling these structures and to compare its performance against Panua-PARDISO 8.2, a state-of-the-art solver.
@@
    Factorization and Selected Inversion: Cholesky factorization is performed, followed by selected inversion using sTiles and Panua-PARDISO 8.2. For each specific test case (a given matrix on a given number of cores), a single execution was performed to measure the runtime of the selected inverse.
@@
    Parallel Scalability Evaluation: To assess parallel performance, these execution times were recorded across a wide range of core counts (1, 2, 4, 8, 16, 32, and 52 cores, the maximum available on our test system). This allows us to observe how each solver’s performance scales with increasing parallelism.
@@
    Performance Analysis and Visualization: The results are analyzed in terms of absolute runtime and relative speedup. In our summary discussions, we sometimes refer to the “best execution time,” which denotes the optimal performance achieved by a solver for a given problem across all tested core counts. This is a key part of the analysis, as parallel overhead can sometimes cause performance to degrade with a higher number of cores, and identifying this optimal point is important.
@@
Figure 7 demonstrates the performance advantage of sTiles over Panua-PARDISO across different matrix IDs ordered by increasing sizes. The performance gap becomes more pronounced as the matrix size increases (IDs 7-18), with sTiles leveraging dense tiles to mitigate computational overhead and memory bandwidth limitations. In some cases, sTiles achieves up to 13.49× speedup compared to Panua-PARDISO.
IV-C Parallel Scalability Evaluation
Figure 8 presents the execution times of sTiles and Panua-PARDISO across different core counts for two representative matrix sizes. For the smaller 10K matrix, both solvers exhibit strong parallel scalability. However, Panua-PARDISO consistently achieves shorter execution times in this scenario. The trend reverses for the larger 500K matrix, where sTiles demonstrates superior parallel scalability and achieves increasingly lower execution times as the core count grows. This performance advantage highlights the efficiency of sTiles’s parallelization strategy on large-scale problems.
Figure 8: Scalability of sTiles and Panua-PARDISO for selected matrix configurations. From left to right: a medium-sized matrix with low bandwidth (ID 7), a similar matrix with a thicker arrowhead structure (ID 10), and a matrix with high bandwidth and computational intensity (ID 12). sTiles demonstrates superior scalability, particularly on larger and more computationally demanding problems.
@@ IV-D Impact of Sparsity on Performance
The impact of sparsity on the computational performance of selected inversion is analyzed using a set of structured matrices with varying densities. Table III summarizes the properties of these matrices, including size, bandwidth, arrowhead thickness, and density. The density values, which exclude the arrowhead portion (see Figure 6), range from 0.010% to 4.101%, providing a broad spectrum of sparsity levels for evaluation.
Figure 9 presents the inverse computation times for both Panua-PARDISO and sTiles solvers across matrices with increasing density. The results illustrate a clear distinction in solver behavior depending on the sparsity characteristics. For matrices with very low density (below 0.1%), Panua-PARDISO exhibits a faster performance than sTiles, primarily due to its optimized multifrontal structure for handling highly sparse matrices. However, as the density increases beyond this threshold, the performance of sTiles stabilizes, while Panua-PARDISO experiences significant computational overhead.
The results indicate that sTiles maintains consistent computational times across a wide range of density values, demonstrating its robustness for handling moderately sparse to dense matrices. In contrast, Panua-PARDISO’s runtime increases substantially with density, reflecting the growing complexity of fill-in and symbolic factorization in direct solvers. For high-density matrices (greater than 1%), sTiles outperforms Panua-PARDISO, making it a more scalable approach for problems with increasing density.
@@
TABLE III: Matrix properties used in Cholesky factorization and selected inversion experiments for sTiles – Set 2. Density values exclude the arrowhead part.
@@
Figure 9: Selected inverse computation times of sTiles and Panua-PARDISO across varying matrix densities. The top plot illustrates results for matrices with a bandwidth of 1500, while the bottom plot corresponds to matrices with a bandwidth of 3000. The x-axis represents matrix IDs, with density values displayed beneath each ID, while the y-axis uses a logarithmic scale to capture the variations in computation time.
IV-E Framework Robustness
While our primary focus is on selected inversion, the underlying sTiles framework is also highly efficient for full matrix inversion. To demonstrate this robustness, we evaluate the performance of sTiles against PLASMA [19, 21], a state-of-the-art numerical library, for full matrix inversion. The comparison highlights the fundamental difference in scheduling strategies: sTiles employs a static scheduling approach designed to maximize data locality and minimize runtime overhead, whereas PLASMA utilizes a dynamic scheduler.
To ensure a fair comparison focused purely on the core computational kernels, our measurements for PLASMA exclusively time the inverse function. The initial data translation to a tile layout (plasma_omp_dtr2desc) and the final translation back to a standard LAPACK layout (plasma_omp_ddesc2tr) are excluded from the timing. This methodology isolates the performance of the inversion algorithm itself from data layout conversion costs.
Figures 10 and 11 present the execution times for both libraries on matrix ID 5 (from Table II) across a range of tile sizes and core counts. The results reveal distinct performance characteristics:
  • Scalability: sTiles demonstrates strong and consistent scalability, with execution time decreasing steadily up to 52 cores. In contrast, PLASMA’s performance stagnates or degrades beyond 16 cores, particularly for smaller tile sizes (40 and 80). This suggests that the overhead from its dynamic scheduler and associated thread contention becomes a significant bottleneck at high core counts.
  • Tile Size Sensitivity: sTiles exhibits lower sensitivity to tile size, achieving robust performance across various configurations without requiring extensive tuning. PLASMA’s performance is more variable with tile size, underscoring the challenge of managing its dynamic task dependencies effectively.
@@ IV-E Framework Robustness
To distill these findings, Figure 12 provides a direct comparison using the optimal tile size for each library (280 for PLASMA and 320 for sTiles). The results illustrate the classic trade-off between static and dynamic scheduling. sTiles is significantly faster at lower core counts (1-16 cores) due to lower overhead. At 32 cores, a crossover occurs where PLASMA’s dynamic scheduler achieves a marginal advantage, likely by better balancing a slightly uneven workload.
However, as the core count increases to 52, the low overhead of sTiles’s static approach proves superior, and it reclaims the performance lead. The flattening curve for PLASMA indicates that the cost of dynamic task management becomes the primary bottleneck, limiting its scalability. This confirms that the static scheduling and data locality optimizations in sTiles are more effective for achieving high parallel efficiency on many-core architectures.
Figure 12: Full inversion comparison using the best tile size for each library (PLASMA: 280, sTiles: 320). sTiles demonstrates superior scalability at high core counts, outperforming PLASMA where parallel efficiency is most critical.
@@ IV-F Accelerating Selected Inversion on GPU
To evaluate the performance of our algorithm on massively parallel hardware, we leveraged a hybrid CPU-GPU implementation. In this model, CPU host threads orchestrate the task dependency graph and enqueue numerically intensive kernels onto an NVIDIA A100 GPU. Each thread manages a dedicated CUDA stream, maximizing concurrency. This subsection analyzes the resulting performance and explores the critical trade-offs involved in GPU acceleration.
IV-F1 Execution Model and Performance Evaluation
Before computation, the entire matrix is transferred to GPU memory. CPU threads then enqueue cuBLAS/cuSOLVER kernels onto their respective streams. Synchronization is handled on the CPU via the shared core_progress array, with host threads waiting until dependencies are met before launching the next GPU kernel.
To demonstrate the potential of this approach, we selected a computationally demanding problem: the inversion of a large matrix (size 50,010, bandwidth 15,000) that has sufficient computational work to saturate the GPU and fits entirely within its 80 GB memory, thereby avoiding out-of-core complexities. For this direct comparison, all experiments were run on the GPU Node (AMD EPYC 7713 vs. NVIDIA A100).
A key factor in performance is the tile size, which was tuned for each architecture: a smaller size of 120 for the CPU to expose task parallelism across many cores, and a much larger size of 600 for the GPU to ensure that each kernel performs enough work to be efficient. This choice creates a fundamental trade-off: the large GPU tile size maximizes kernel throughput but reduces the overall number of parallel tasks available in the algorithm.
TABLE IV: CPU vs GPU performance comparison for the two-phase selected inversion.
Cores/ Streams CPU Time (s) GPU Time (s) GPU Time + Data (s)
1 533.062 5.020 7.044
2 282.714 2.517 4.524
4 142.883 2.416 4.366
8 75.225 2.015 3.954
16 39.575 2.285 4.111
32 26.861 1.913 3.562
64 20.382 2.351 3.949
[…]
The observed speedup, while significant, is below the theoretical peak performance ratio of the hardware. This is an expected outcome due to three main factors:
  1. 1.
    Reduced Task Parallelism: The large tile size required for GPU efficiency limits the number of concurrent tasks available in the algorithm’s dependency graph.
  2. 2.
    Synchronization Overhead: The dependency checks are managed by CPU threads, which introduces latency.
  3. 3.
    Data Transfer Cost: The reported “GPU Time + Data” in Table IV includes only the final transfer of the inverted result from device to host memory. The initial matrix and Cholesky tiles are already on the GPU, and the inverse is initialized to zero directly in device memory, so no additional transfers occur during the selected inversion phase.
In conclusion, our hybrid CPU-GPU model provides a substantial performance benefit for large-scale structured matrix inversion. The results validate that the sTiles framework can effectively exploit GPU acceleration, with performance shaped by a clear trade-off between single-kernel efficiency (large tiles) and overall algorithmic parallelism (small tiles).